# CAML: Collaborative Auxiliary Modality Learning for Multi-Agent Systems

### v1

@@ Abstract
Multi-modality learning has become a crucial technique for improving the performance of machine learning applications across domains such as autonomous driving, robotics, and perception systems. While existing frameworks such as Auxiliary Modality Learning (AML) effectively utilize multiple data sources during training and enable inference with reduced modalities, they primarily operate in a single-agent context. This limitation is particularly critical in dynamic environments, such as connected autonomous vehicles (CAV), where incomplete data coverage can lead to decision-making blind spots. To address these challenges, we propose Collaborative Auxiliary Modality Learning (CAML), a novel multi-agent multi-modality framework that enables agents to collaborate and share multimodal data during training while allowing inference with reduced modalities per agent during testing. We systematically analyze the effectiveness of CAML from the perspective of uncertainty reduction and data coverage, providing theoretical insights into its advantages over AML. Experimental results in collaborative decision-making for CAV in accident-prone scenarios demonstrate that CAML achieves up to a ${\bf 58.13}\%$ improvement in accident detection. Additionally, we validate CAML on real-world aerial-ground robot data for collaborative semantic segmentation, achieving up to a ${\bf 10.61}\%$ improvement in mIoU.
Figure 1: Illustration of CAML. CAML enables multiple agents to collaborate and share multimodal data during training while allowing for runtime inference with reduced modalities per agent during testing. Additionally, the number of agents can vary between training and testing, ensuring flexibility and robustness in deployment.
@@ 1 Introduction
Multi-modality learning has become an essential approach in a wide range of machine learning applications, particularly in areas such as autonomous driving (El Madawi et al., 2019; Xiao et al., 2020; Gao et al., 2018) , robotics (Noda et al., 2014; Lee et al., 2020), and perception systems (Zhuang et al., 2021; Bayoudh et al., 2022), where the availability of multiple data sources (e.g., RGB images, LiDAR, radar, etc.) improves model performance by providing complementary information. However, these multi-modality systems often suffer from increased computational complexity and latency at inference time. Moreover, some modalities may not be consistently available or reliable in real-world conditions, necessitating strategies that can compensate for missing modalities during inference.
@@
Recent work on machine learning (Hoffman et al., 2016; Wang et al., 2018; Garcia et al., 2018, 2019; Piasco et al., 2021) aims to address these problems by allowing models to leverage additional modalities during training while enabling inference using fewer or even a single modality. For example, a model might be trained using both RGB and LiDAR data, but during deployment, it only requires RGB data to operate. These approaches reduce the computational burden and accommodates real-world conditions where certain sensors may be unavailable. Shen et al. (2023) formalized these learning tasks as Auxiliary Modality Learning (AML). The AML framework successfully reduces the dependency on expensive or unreliable modalities, but it focuses on the single-agent setting, where an individual model is trained to handle reduced modalities during inference.
@@
Despite the benefits of AML, several gaps remain. First, a major limitation in the current AML framework is the inability to exploit collaboration between agents, particularly in dynamic environments such as connected autonomous vehicles (CAV). In such scenarios, data coverage from a single agent is often incomplete because of occlusion or limited sensor range, leading to blind spots or increased uncertainty in decision-making. Second, the information from multiple modalities can complement each other across agents, especially in multi-agent settings such as vehicle-to-vehicle (V2V) communication or collaborative robotics. Different agents may have access to complementary sensory information, which could be shared to have agents make more informed and safer decisions, notably in accident-prone scenarios. However, current AML approaches do not exploit this potential for collaboration.
@@
To bridge these gaps, we propose Collaborative Auxiliary Modality Learning (CAML), a novel framework for multi-agent multi-modality systems that allows agents to collaborate and share multimodal data during training, but enables inference with reduced modalities per agent during testing, as illustrated in Figure 1. CAML leverages knowledge distillation (Hinton, 2015), transferring knowledge from a teacher model into a student model. This enables the student to operate with missing modalities during inference. For instance, in autonomous driving, multiple vehicles can share sensor information such as LiDAR and RGB images during training to build more robust representations, while during runtime testing, each vehicle performs inference using only RGB images.
CAML addresses two key challenges: (1) It reduces uncertainty and enhances data coverage in dynamic environments by leveraging complementary information from multiple agents. (2) It maintains efficient, modality-reduced inference during testing. Unlike previous work that either focuses on multi-agent collaboration but without addressing modality reduction at test time, or tackles multi-modality learning in single-agent settings, CAML unifies these concepts. Through collaboration, CAML enables agents to compensate for each other’s blind spots, resulting in more informed prediction or decision-making even when some modalities are unavailable at deployment. In summary, our work offers the following key contributions:
  • We introduce CAML, a novel framework for multi-agent systems that allows agents to share multimodal data during training, while performing efficient, reduced-modality inference during testing. By leveraging the strengths of multi-agent collaboration, CAML can reduce estimation uncertainty and integrate complementary information, capturing a broader and more detailed data representation.
  • We systematically analyze the effectiveness of CAML from the perspective of uncertainty reduction and enhanced data coverage, providing theoretical insights into its advantages over AML.
@@
  • We validate CAML through experiments in collaborative decision-making for connected autonomous driving in accident-prone scenarios, and collaborative semantic segmentation for real-world data of aerial-ground robots. CAML achieves up to ${\bf 58.13}\%$ improvement in accident detection for autonomous driving, and up to ${\bf 10.61}\%$ improvement for more accurate semantic segmentation.
@@ 2 Related Work
Collaboration in multi-agent systems has been widely studied across fields such as autonomous driving and robotics. In autonomous driving, prior research has explored various strategies, including spatio-temporal graph neural networks (Gao et al., 2024), LiDAR-based end-to-end systems (Cui et al., 2022), decentralized cooperative lane-changing (Nie et al., 2016) and game-theoretic models (Hang et al., 2021). In robotics, Mandi et al. (2024) presented a hierarchical multi-robot collaboration approach using large language models, while Zhou et al. (2022) proposed a perception framework for multi-robot systems built on graph neural networks. A review of multi-robot systems in search and rescue operations was provided by (Queralta et al., 2020), and Bae et al. (2019) developed a reinforcement learning (RL) method for multi-robot path planning. Additionally, various communication mechanisms, such as Who2com (Liu et al., 2020b), When2com (Liu et al., 2020a), and Where2comm (Hu et al., 2022), have been created to optimize agent interactions.
Despite these advancements, existing multi-agent collaboration frameworks remain limited by their focus on specific tasks and the assumption that agents will have consistent access to the same data modalities during both training and testing, an assumption that may not hold in real-world applications. To address these gaps, our framework, CAML, enables agents to collaborate during training by sharing multimodal data, but at test time, each agent performs inference using reduced modality. This reduces the dependency on certain modalities for deployment, while still allowing agents to leverage additional data during training to enhance overall performance and robustness.
@@
Auxiliary Modality Learning (AML) (Shen et al., 2023) has emerged as an effective solution to reduce computational costs and the amount of input data required for inference. By utilizing auxiliary modalities during training, AML minimizes reliance on those modalities at inference time. For example, Hoffman et al. (2016) introduced a method that incorporates depth images during training to enhance test-time RGB-only detection models. Similarly, Wang et al. (2018) proposed PM-GANs to learn a full-modal representation using data from partial modalities, while Garcia et al. (2018, 2019) developed approaches that use depth and RGB videos during training but rely solely on RGB data for testing. Piasco et al. (2021) created a localization system that predicts depth maps from RGB query images at test time. Building on these works, Shen et al. (2023) formalized the AML framework, systematically classifying auxiliary modality types and AML architectures.
However, existing AML frameworks are typically designed for single-agent settings, failing to exploit the potential benefits of multi-agent collaboration for improving multimodal learning. CAML allows agents to collaboratively learn richer multimodal representations during training. This approach mitigates the loss of information when modalities are reduced during inference, as the learned features are reinforced by data shared across agents.
@@
Knowledge distillation (KD) (Hinton, 2015) is a widely used technique in many domains to reduce computation by transferring knowledge from a large, complex model (teacher) to a simpler model (student). In computer vision, Gou et al. (2021) provided a comprehensive survey of KD applications, while Beyer et al. (2022) conducted an empirical investigation to develop a robust and effective recipe for making State-of-the-Art (SOTA) large-scale models more practical. Additionally, Tung & Mori (2019) introduced a KD loss function that aligns the training of a student network with input pairs producing similar activation in the teacher network. In natural language processing, Xu et al. (2024) reviewed the applications of KD in LLMs, while Sun et al. (2019) proposed a Patient KD method to compress larger models into lightweight counterparts that maintain effectiveness. Hahn & Choi (2019) also suggested a KD approach that leverages the soft target probabilities of the training model to train other neural networks. In autonomous driving, Lan & Tian (2022) presented an approach for visual detection, Cho et al. (2023); Sautier et al. (2022) used KD for 3D object detection.
Notice that existing KD mostly distills knowledge from a larger model to a smaller one to reduce computation, Shen et al. (2023) aimed to design a cross-modality learning approach using KD to utilize the hidden information from auxiliary modalities within the AML framework. But AML is limited by the scope of a single-agent paradigm, missing opportunities for collaborative knowledge sharing across agents. In contrast, we leverage KD within multi-agent settings, where the teacher models are trained with access to shared multimodal data (e.g., RGB and LiDAR) from multiple agents. By distilling this collaborative knowledge into each agent’s reduced modality (e.g., RGB), CAML enables robust inference during deployment, even with fewer modalities. This collaborative distillation process enhances each agent’s performance by providing richer, complementary knowledge from the collaborative training phase.
@@ 3 Collaborative Auxiliary Modality Learning
In AML (Shen et al., 2023), which operates in a single-agent framework, the missing modalities during testing are referred to as auxiliary modalities, while those that remain available are called the main modality. In contrast, in our framework CAML, each agent can process a different number of modalities during training and different agents can have different main modalities and auxiliary modalities. There is no correlation between the number of agents and the number of modalities.
@@
We define our problem in both training and testing phases. In the training phase, we consider a multi-agent system with $N$ agents collaboratively completing a task. The set of agents is denoted as ${\mathcal{A}}_{train}=\{{\mathcal{A}}_{1},{\mathcal{A}}_{2},\ldots,{\mathcal{A}}_{N}\}$. The observations of all agents are denoted as $X=\{x_{1},x_{2},\ldots,x_{N}\}$, where $x_{i}$ is the observation acquired by the $i$-th agent ${\mathcal{A}}_{i}\in{\mathcal{A}}_{train}$. The ground truth label is denoted as $Y$, which can be an object label, semantic class, or a control command (e.g., brake for an autonomous vehicle). The set of modalities is denoted as ${\mathcal{I}}_{train}=\{{\mathcal{I}}_{1},{\mathcal{I}}_{2},\ldots,{\mathcal{I}}_{K}\}$, such as RGB, LiDAR, Depth, etc, where $K$ is the number of modalities avaiable during training. During training, each agent has access to all these $K$ modalities. In the testing phase, we assume there are $M$ agents. The set of test agents is denoted as ${\mathcal{A}}_{test}=\{{\mathcal{A}}_{1},{\mathcal{A}}_{2},\ldots,{\mathcal{A}}_{M}\}$. In addition, the set of modalities is denoted as ${\mathcal{I}}_{test}$, which is a subset of ${\mathcal{I}}_{train}$. The number of modalities available during testing is denoted as $L$, where $L\leq K$. The set of agents that have access to the $j$-th modality ${\mathcal{I}}_{j}\in{\mathcal{I}}_{test}$ is denoted as ${\mathcal{A}}_{test}^{{\mathcal{I}}_{j}}$, where ${\mathcal{A}}_{test}^{{\mathcal{I}}_{j}}\in{\mathcal{A}}_{test}$, and the number of agents in this set is given by $|{\mathcal{A}}_{test}^{{\mathcal{I}}_{j}}|=M_{j}$. This means that during testing, each agent may have access to different number of modalities.
@@
Given the problem definition, we aim to estimate the posterior distribution $P(y|X)$ of the ground truth label $y$ given all agents’ observations $X$. During training, we train both a teacher model where each agent has access to all modalities in ${\mathcal{I}}_{train}$ and a student model where each agent has access to partial modalities in ${\mathcal{I}}_{test}$. We employ Knowledge Distillation (KD) to transfer the knowledge derived from the teacher model to the student model, enabling the student to benefit from additional information, as illustrated in Figure 2. At test time, we perform inference using the student model, which relies on the test modality observations $X^{test}$.

### v2

@@ Abstract
Multi-modal learning has become a crucial technique for improving the performance of machine learning applications across domains such as autonomous driving, robotics, and perception systems. However, in certain scenarios, particularly in resource-constrained environments, some modalities available during training may be absent during inference. While existing frameworks effectively utilize multiple data sources during training and enable inference with reduced modalities, they are primarily designed for single-agent settings. This poses a critical limitation in dynamic environments such as connected autonomous vehicles (CAV), where incomplete data coverage can lead to decision-making blind spots. Conversely, some works explore multi-agent collaboration but without addressing missing modality at test time. To overcome these limitations, we propose Collaborative Auxiliary Modality Learning (CAML), a novel multi-modal multi-agent framework that enables agents to collaborate and share multi-modal data during training, while allowing inference with reduced modalities during testing. Experimental results in collaborative decision-making for CAV in accident-prone scenarios demonstrate that CAML achieves up to a ${\bf 58.1}\%$ improvement in accident detection. Additionally, we validate CAML on real-world aerial-ground robot data for collaborative semantic segmentation, achieving up to a ${\bf 10.6}\%$ improvement in mIoU.
@@ 1 Introduction
Multi-modal learning has become an essential approach in a wide range of machine learning applications, particularly in areas such as autonomous driving [6, 37, 7], robotics [27, 20], and perception systems [42, 2], where the availability of multiple data sources (e.g., RGB images, LiDAR, radar) improves model performance by providing complementary information. However, these multi-modal systems often incur increased computational overhead and latency at inference time. More critically, in real-world conditions, some modalities may be unavailable, or too costly to acquire at test time, which challenges the reliability and efficiency of such systems.
@@
Recent work in machine learning [16, 36, 9, 10, 28] has sought to address these problems by enabling models to leverage additional modalities during training while supporting inference with fewer modalities. Such approaches reduce computational costs and improve robustness in conditions where some sensors may be unavailable at inference time. Shen et al. [32] formalized these tasks under the framework of Auxiliary Modality Learning (AML), which effectively reduces the dependency on expensive or unreliable modalities.
@@
Despite the benefits of AML, notable limitations remain. AML is primarily designed for single-agent settings, where an individual model is trained to handle reduced modalities during inference. A key challenge in current AML frameworks is insufficient data coverage, particularly in dynamic environments such as connected autonomous vehicles (CAV). In these scenarios, a single agent’s data is often incomplete due to occlusions or limited sensor range, resulting in blind spots and increased uncertainty in decision-making. In contrast, multi-agent systems, enabled by technologies such as vehicle-to-vehicle (V2V) communication or collaborative robotics, can access complementary sensory information from other agents. Sharing this information allows for more informed and safer decisions, especially in accident-prone scenarios. However, existing single-agent AML approaches fail to leverage this collaborative potential.
Figure 1: Illustration of CAML. CAML enables (1) multiple agents to collaborate and share multi-modal data during training while allowing for runtime inference with reduced modalities during testing; (2) the number of agents can vary between training and testing, ensuring flexibility and robustness in deployment.
@@
To address these limitations, we propose Collaborative Auxiliary Modality Learning (CAML), a novel framework for multi-modal, multi-agent systems. CAML enables agents to collaborate and share multi-modal data during training, while supporting inference with reduced modalities at test time, as illustrated in Fig. 1. This collaborative learning process enhances scene understanding and data coverage by allowing agents to compensate for each other’s blind spots, leading to more informed predictions and decision-making. To ensure effective deployment in scenarios where only partial modalities are available (e.g., due to sensor failures or resource constraints), CAML utilizes knowledge distillation [15] to transfer knowledge from a teacher model, trained with full-modality inputs, to a student model that learns to perform effectively with reduced modalities at inference time. This distillation process embeds full-modality knowledge into the student model, preserving the benefits of collaborative training and ensuring robust inference even when some modalities are absent. For instance, in autonomous driving, multiple vehicles can share sensor information such as LiDAR and RGB images during training to build more robust representations, while during runtime testing, each vehicle performs inference using only RGB images.
Unlike prior work that either focuses on multi-agent collaboration but without addressing missing modality at test time, or explores AML ideas solely in single-agent settings, CAML unifies these concepts. Overall, this work offers the following key contributions:
  • We introduce CAML, a novel framework for multi-agent systems that allows agents to share multi-modal data during training, while supporting reduced-modality inference during testing. This collaborative approach enhances data coverage, improves robustness to missing modalities, which is crucial for deployment in in dynamic, resource-constrained environments.
@@
  • We validate CAML through experiments in collaborative decision-making for connected autonomous driving in accident-prone scenarios, and collaborative semantic segmentation for real-world data of aerial-ground robots. CAML achieves up to ${\bf 58.1}\%$ improvement in accident detection for autonomous driving, and up to ${\bf 10.6}\%$ improvement for more accurate semantic segmentation.
@@ 2 Related Work
Collaboration in multi-agent systems has been widely studied across fields such as autonomous driving and robotics. In autonomous driving, prior research has explored various strategies, including spatio-temporal graph neural networks [8], LiDAR-based end-to-end systems [5], decentralized cooperative lane-changing [26] and game-theoretic models [13]. In robotics, Mandi et al. [25] presented a hierarchical multi-robot collaboration approach using large language models, while Zhou et al. [40] proposed a perception framework for multi-robot systems built on graph neural networks. A review of multi-robot systems in search and rescue operations was provided by [30], and Bae et al. [1] developed a reinforcement learning (RL) method for multi-robot path planning. Additionally, various communication mechanisms, such as Who2com [22], When2com [21], and Where2comm [17], have been created to optimize agent interactions.
Despite these advancements, existing multi-agent collaboration frameworks remain limited by their focus on specific tasks and the assumption that agents will have consistent access to the same data modalities during both training and testing, an assumption that may not hold in real-world applications. To address these gaps, our framework CAML enables agents to collaborate during training by sharing multi-modal data, but at test time, each agent performs inference using reduced modality. This reduces the dependency on certain modalities for deployment, while still allowing agents to leverage additional information during training to enhance overall performance and robustness.
@@
Auxiliary Modality Learning (AML) [32] has emerged as an effective solution to reduce computational costs and the amount of input data required for inference. By utilizing auxiliary modalities during training, AML minimizes reliance on those modalities at inference time. For example, Hoffman et al. [16] introduced a method that incorporates depth images during training to enhance test-time RGB-only detection models. Similarly, Wang et al. [36] proposed PM-GANs to learn a full-modal representation using data from partial modalities, while Garcia et al. [9, 10] developed approaches that use depth and RGB videos during training but rely solely on RGB data for testing. Piasco et al. [28] created a localization system that predicts depth maps from RGB query images at test time. Building on these works, Shen et al. [32] formalized the AML framework, systematically classifying auxiliary modality types and AML architectures.
However, existing AML frameworks are typically designed for single-agent settings, failing to exploit the potential benefits of multi-agent collaboration for improving multi-modal learning. In contrast, CAML allows agents to collaboratively learn richer multi-modal representations. By sharing complementary information, this approach enhances overall data coverage and mitigates information loss during reduced-modality inference, which is particularly important when an individual agent’s observations are incomplete due to occlusions or limited sensor range.
@@
Knowledge distillation (KD) [15] is a widely used technique in many domains to reduce computation by transferring knowledge from a large, complex model (teacher) to a simpler model (student). In computer vision, Gou et al. [11] provided a comprehensive survey of KD applications, while Beyer et al. [3] conducted an empirical investigation to develop a robust and effective recipe for making large-scale models more practical. Additionally, Tung and Mori [35] introduced a KD loss function that aligns the training of a student network with input pairs producing similar activation in the teacher network. In natural language processing, Xu et al. [38] reviewed the applications of KD in LLMs, while Sun et al. [33] proposed a Patient KD method to compress larger models into lightweight counterparts that maintain effectiveness. Hahn and Choi [12] also suggested a KD approach that leverages the soft target probabilities of the training model to train other neural networks. In autonomous driving, Lan and Tian [19] presented an approach for visual detection, Cho et al. [4], Sautier et al. [31] used KD for 3D object detection.
Notice that existing KD mostly distills knowledge from a larger model to a smaller one to reduce computation, Shen et al. [32] aimed to design a cross-modal learning approach using KD to utilize the hidden information from auxiliary modalities within the AML framework. But AML is limited by the scope of a single-agent paradigm. In contrast, we leverage KD within multi-agent settings, where the teacher models are trained with access to shared multi-modal data (e.g., RGB and LiDAR) from multiple agents. By distilling this collaborative knowledge into each agent’s reduced modality (e.g., RGB), CAML enables robust inference during deployment, even with fewer modalities. This collaborative distillation process enhances each agent’s performance by providing richer, complementary knowledge from the collaborative training phase.
@@ 3 Collaborative Auxiliary Modality Learning
In single-agent frameworks such as AML [32], the missing modalities during testing are referred to as auxiliary modalities, while those that remain available are called the main modality. In contrast, in our framework CAML, each agent can process a different number of modalities during training and different agents can have different main modalities and auxiliary modalities. There is no correlation between the number of agents and the number of modalities.
Figure 2: Overview of the CAML Approach Pipeline. The teacher model (top) aggregates and shares multi-modal embeddings across agents to make predictions using the full set of modalities. In contrast, the student model (bottom) processes a subset of modalities per agent and shares them to form a multi-modal embedding. Through knowledge distillation from the teacher, the student learns to produce robust predictions despite missing modalities, enabling effective inference during deployment. In the teacher model, the set of agents is denoted as ${\mathcal{A}}_{\text{train}}=\{{\mathcal{A}}_{1},{\mathcal{A}}_{2},\ldots,{\mathcal{A}}_{N}\}$. The set of modalities is denoted as ${\mathcal{I}}_{\text{train}}$. The observations of all agents are denoted as $X=\{x_{1},x_{2},\ldots,x_{N}\}$, where $x_{i}^{k}$ is the observation acquired by the $i$-th agent ${\mathcal{A}}_{i}\in{\mathcal{A}}_{\text{train}}$ for the $k$-th modality. In the student model, the set of agents is denoted as ${\mathcal{A}}_{\text{test}}=\{{\mathcal{A}}_{1},{\mathcal{A}}_{2},\ldots,{\mathcal{A}}_{M}\}$. The set of modalities is denoted as ${\mathcal{I}}_{\text{test}}$, which is a subset of ${\mathcal{I}}_{\text{train}}$. The set of agents that have access to the $j$-th modality ${\mathcal{I}}_{j}\in{\mathcal{I}}_{\text{test}}$ is denoted as ${\mathcal{A}}_{\text{test}}^{{\mathcal{I}}_{j}}$, and the number of agents in this set is given by $|{\mathcal{A}}_{\text{test}}^{{\mathcal{I}}_{j}}|=M_{j}$, with $x_{M_{j}}^{j}$ represents the observation acquired by the $M_{j}$-th agent ${\mathcal{A}}_{M_{j}}\in{\mathcal{A}}_{\text{test}}$ for the $j$-th modality.
@@
We define our problem in both training and testing phases. In the training phase, we consider a multi-agent system with $N$ agents collaboratively completing a task. The set of agents is denoted as ${\mathcal{A}}_{\text{train}}=\{{\mathcal{A}}_{1},{\mathcal{A}}_{2},\ldots,{\mathcal{A}}_{N}\}$. The observations of all agents are denoted as $X=\{x_{1},x_{2},\ldots,x_{N}\}$, where $x_{i}$ is the observation acquired by the $i$-th agent ${\mathcal{A}}_{i}\in{\mathcal{A}}_{\text{train}}$. The ground truth is denoted as $Y$, which can be an object label, semantic class, or a control command (e.g., brake for an autonomous vehicle). The set of modalities is denoted as ${\mathcal{I}}_{\text{train}}=\{{\mathcal{I}}_{1},{\mathcal{I}}_{2},\ldots,{\mathcal{I}}_{K}\}$, such as RGB, LiDAR, Depth, etc, where $K$ is the number of modalities available during training. During training, each agent has access to all these $K$ modalities. In the testing phase, we assume there are $M$ agents. The set of test agents is denoted as ${\mathcal{A}}_{\text{test}}=\{{\mathcal{A}}_{1},{\mathcal{A}}_{2},\ldots,{\mathcal{A}}_{M}\}$. In addition, the set of modalities is denoted as ${\mathcal{I}}_{\text{test}}$, which is a subset of ${\mathcal{I}}_{\text{train}}$. The number of modalities available during testing is denoted as $L$, where $L\leq K$. The set of agents that have access to the $j$-th modality ${\mathcal{I}}_{j}\in{\mathcal{I}}_{\text{test}}$ is denoted as ${\mathcal{A}}_{\text{test}}^{{\mathcal{I}}_{j}}$, where ${\mathcal{A}}_{\text{test}}^{{\mathcal{I}}_{j}}\in{\mathcal{A}}_{\text{test}}$, and the number of agents in this set is given by $|{\mathcal{A}}_{\text{test}}^{{\mathcal{I}}_{j}}|=M_{j}$. This means that during testing, each agent may have access to different number of modalities.
@@
Given the problem definition, we aim to estimate the posterior distribution $P(y|X)$ of the ground truth label $y$ given all agents’ observations $X$. During training, we first train a teacher model where each agent has access to all modalities in ${\mathcal{I}}_{\text{train}}$ and then a student model where each agent has access to partial modalities in ${\mathcal{I}}_{\text{test}}$. We employ Knowledge Distillation (KD) to transfer the richer multi-modal knowledge from the teacher to the student, enabling the student to benefit from information beyond its own input, as illustrated in Fig. 2. At test time, we perform inference using the student model, which operates solely on the available test modality observations $X_{\text{test}}$.